# Challenge-Response Quantum Reinforcement Learning with Application to Quantum-Assisted Authentication

### v1

@@ Abstract
Quantum reinforcement learning (QRL) has emerged as a promising research direction that integrates quantum information processing into reinforcement learning frameworks. While many existing QRL studies apply quantum agents to classical environments, it has been realized that the potential advantages of QRL are most naturally explored in environments that exhibit intrinsically quantum characteristics, where the agent’s observations and interactions arise from quantum processes. In this work, we propose a quantum reinforcement learning environment formulated as a challenge-response task with hidden information. In the proposed environment, Alice encodes a classical bit into the parameters of a quantum circuit, while Bob, with a trained reinforcement learning agent, interacts with a limited number of quantum state copies to infer the hidden bit. The agent must select measurement strategies and decide when to terminate the interaction under explicit resource constraints. To study the solvability of the proposed environment, we consider three agents: a purely classical agent, a lightweight hybrid agent and a deep hybrid agent. Through experiments, we analyze the trade-off between inference accuracy and quantum resource consumption under varying interaction penalties. Our results show that the lightweight hybrid agent achieves reliable inference using as few as two quantum state copies, outperforming both the classical baseline and the deep hybrid agent in highly resource-constrained regimes. We further evaluate robustness under realistic quantum noise models and discuss the relevance of the proposed environment for security-oriented applications, including quantum-assisted authentication.
@@ I Introduction
In response to these challenges, several works have investigated QRL performance in a different RL environments. For example, QRL has been applied to solve maze-based navigation environments, which are often used as simplified testbeds for analyzing learning behavior [8, 29]. In [13], hybrid quantum agents were applied to high-dimensional RL environments such as Atari games, where the proposed hybrid model was capable of solving Pong and achieving performance comparable to a classical baseline in Breakout. QRL has also been explored as a modeling framework in cognitive science, such as in [20], where QRL-based models demonstrated competitive performance compared to classical RL models in explaining behavioral data. Quantum-inspired and quantum-enhanced reinforcement learning approaches have also been explored in healthcare decision-making, where quantum optimization techniques are used to improve policy learning for personalized treatment planning [28]. In addition, quantum-inspired reinforcement learning has been investigated in classical environments modeling complex real-world decision problems, such as secure and sustainable supply chain systems [9]
@@ I-B Organization
The remainder of this paper is organized as follows. Section II provides background on reinforcement learning and quantum reinforcement learning. Section III presents the quantum challenge–response circuit underlying the proposed environment, followed by a detailed description of the environment in Section IV. Section V describes our agents. Section VII introduces the evaluation metrics, and the experimental results are present in Section VIII. Section IX discusses a potential application of the proposed environment, and Section X concludes the paper.
@@ V Agent Architectures
The agent is trained to interact with the environment in order to infer the hidden classical bit based solely on the measurement outcomes it observes. While the measurement probabilities are governed by the mathematical expressions in (6), these relationships are not provided explicitly to the agent. Instead, the agent learns an effective probing and stopping strategy through RL by observing the outcomes of its interactions. The agent learns to select test angles that yield informative measurement statistics and to decide when sufficient evidence has been collected to make a final guess. This requires implicitly capturing the relationship between probe angles and measurement outcomes for the two possible hidden bits, rather than explicitly minimizing the difference between $\phi_{1}$ and $\phi_{2}$.
@@ V-B D-agent
In this agent, the actor is implemented as a parameterized quantum circuit, while the critic is realized as a classical neural network, as illustrated in Fig. 2. The critic estimates the state value and is used solely as a baseline during training, whereas action selection is entirely governed by the quantum policy. The quantum actor adopts a hardware-efficient ansatz, which is known to provide strong expressivity with shallow circuit depth [30, 15]. The circuit is composed of alternating encoding and variational layers. The encoding layer embeds the observation into the quantum state using single-qubit rotation gates $R_{y}$ and $R_{x}$. The variational layer consists of trainable single-qubit rotations, again using $R_{y}$ and $R_{x}$. followed by entangling controlled-Z (CZ) gates. These alternating layers enable the circuit to learn expressive policies while remaining compatible with near-term quantum hardware constraints.
@@ VI-A Agents
Agent Quantum Depth Classical Depth LR (Quantum Actor) LR (Classical Policy) LR (Critic) Batch Size
C-agent 1 (128 neurons) 0.01 30
D-agent 4 1 (64 neurons) $\{0.01,0.01,0.005\}$ 0.001 30
S-agent 2 1 (128 neurons) 0.01 30
@@ VII Evaluation Metrics
where $b_{i}$ denotes the true hidden bit in episode $i$, $\hat{b}_{i}$ is the agent’s predicted bit at episode, $\mathbb{I}(\cdot)$ is the indicator function, and $N_{\mathrm{ep}}$ is the batch size ($N_{\mathrm{ep}}=30$ in our experiments).
@@ VIII-B Low-Penalty Environment
From these results, we conclude that both D-agent and S-agent outperform the C-agent, as they require fewer quantum copies to make a correct inference while achieving higher accuracy. This demonstrates that hybrid agents are more resource-efficient in the low-penalty setting. This trend is further confirmed by the confusion matrices in Fig. 7, which show that both D-agent and S-agent achieve more balanced and reliable predictions across the two hidden bit values compared to the C-agent.
We note that S-agent employs a simpler quantum circuit than D-agent. In scenarios where circuit complexity is a critical consideration, S-agent offers a better trade-off by achieving accuracy comparable to D-agent while requiring fewer quantum states and reduced circuit complexity, and thus lower implementation cost. Moreover, although S-agent achieves slightly lower accuracy than D-agent in the low-penalty environment, its performance remains comparable while maintaining its advantage under the strict interaction constraints of the high-penalty setting. This consistency highlights S-agent as a robust and resource-efficient policy across varying interaction costs.
@@ X Conclusion
Specifically, the lightweight S-agent consistently achieved reliable inference using as few as two quantum state copies, outperforming both the classical baseline and the deeper quantum policy in highly resource-constrained settings. This finding highlights that increased quantum circuit depth does not necessarily translate to improved performance and that compact hybrid policies can offer a favorable balance between expressivity, stability, and resource efficiency. Robustness evaluations under realistic quantum noise models further showed that the hybrid agents maintain improved inference performance compared to the C-agent across a broad range of noise strengths.
@@
Another potential application for our proposed environment information hiding. In such a setting, classical bits are encoded into quantum states using secret phase angles. The resulting quantum states are stored in quantum memory, while the corresponding angular domain is stored separately as classical side information. Information extraction is only possible for an authorized party that has access to both the stored quantum states and the associated angular domain. This knowledge enables the extractor to train an RL agent tailored to the encoding parameters, which can then interact with the quantum states to reliably recover the hidden bits. Without simultaneous access to both the quantum states and the correct angular domain, any attempt to extract the hidden information yields incorrect results, effectively concealing the embedded data. This dual-dependence on quantum memory and classical side information highlights the potential of RL-based quantum challenge–response mechanisms as a viable tool for information hiding.

### v2

@@ Abstract
Quantum reinforcement learning (QRL) has emerged as a promising research direction that integrates quantum information processing into reinforcement learning frameworks. While many existing QRL studies apply quantum agents to classical environments, it has been realized that the potential advantages of QRL are most naturally explored in environments that exhibit intrinsically quantum characteristics, where the agent’s observations and interactions arise from quantum processes. In this work, we propose a quantum reinforcement learning environment formulated as a challenge-response task with hidden information. In the proposed environment, Alice encodes a classical bit into the parameters of a quantum circuit, while Bob, with a trained reinforcement learning agent, interacts with a limited number of quantum state copies to infer the hidden bit. The agent must select measurement strategies and decide when to terminate the interaction under explicit resource constraints. To study the solvability of the proposed environment, we consider three agents: a purely classical agent, a lightweight hybrid agent and a deep hybrid agent. Through experiments, we analyze the trade-off between inference accuracy and quantum resource consumption under varying interaction penalties. Our results show that the lightweight hybrid agent achieves reliable inference using as few as two quantum state copies, outperforming both the classical baseline and the deep hybrid agent across both high and low-penalty regimes. We further evaluate robustness under realistic quantum noise models and discuss the relevance of the proposed environment for security-oriented applications, including quantum-assisted authentication.
@@ I Introduction
In response to these challenges, several works have investigated QRL performance in a different RL environments. For example, QRL has been applied to solve maze-based navigation environments, which are often used as simplified testbeds for analyzing learning behavior [8, 29]. In [13], hybrid quantum agents were applied to high-dimensional RL environments such as Atari games, where the proposed hybrid model was capable of solving Pong and achieving performance comparable to a classical baseline in Breakout. QRL has also been explored as a modeling framework in cognitive science, such as in [20], where QRL-based models demonstrated competitive performance compared to classical RL models in explaining behavioral data. Quantum-inspired and quantum-enhanced reinforcement learning approaches have also been explored in healthcare decision-making, where quantum optimization techniques are used to improve policy learning for personalized treatment planning [28]. In addition, quantum-inspired reinforcement learning has been investigated in classical environments modeling complex real-world decision problems, such as secure and sustainable supply chain systems [9].
@@ I-B Organization
The remainder of this paper is organized as follows. Section II provides background on reinforcement learning and quantum reinforcement learning. Section III presents the quantum challenge–response circuit underlying the proposed environment, followed by a detailed description of the environment in Section IV. Section V describes our agents. Section VI presents the experimental setup, Section VII introduces the evaluation metrics, and the experimental results are present in Section VIII. Section IX discusses a potential application of the proposed environment, and Section X concludes the paper.
@@ V Agent Architectures
The agent is trained to interact with the environment in order to infer the hidden classical bit based solely on the measurement outcomes it observes. While the measurement probabilities are governed by the mathematical expressions in Eq. 6, these relationships are not provided explicitly to the agent. Instead, the agent learns an effective probing and stopping strategy through RL by observing the outcomes of its interactions. The agent learns to select test angles that yield informative measurement statistics and to decide when sufficient evidence has been collected to make a final guess. This requires implicitly capturing the relationship between probe angles and measurement outcomes for the two possible hidden bits, rather than explicitly minimizing the difference between $\phi_{1}$ and $\phi_{2}$.
@@ V-B D-agent
In this agent, the actor is implemented as a parameterized quantum circuit, while the critic is realized as a classical neural network, as illustrated in Fig. 2. The critic estimates the state value and is used solely as a baseline during training, whereas action selection is entirely governed by the quantum policy. The quantum actor adopts a hardware-efficient ansatz, which is known to provide strong expressivity with shallow circuit depth [30, 15]. The circuit is composed of alternating encoding and variational layers. The encoding layer embeds the observation into the quantum state using single-qubit rotation gates $R_{y}$ and $R_{z}$. The variational layer consists of trainable single-qubit rotations, again using $R_{y}$ and $R_{z}$. followed by entangling controlled-Z (CZ) gates. These alternating layers enable the circuit to learn expressive policies while remaining compatible with near-term quantum hardware constraints.
@@ VI-A Agents
Agent Quantum Depth Classical Depth LR (Quantum Actor) LR (Classical Policy) LR (Critic) Training Batch Eval. Episodes $\boldsymbol{\gamma}$
C-agent 1 (128 neurons) 0.01 30 1000 1
D-agent 4 1 (64 neurons) $\{0.01,\;0.01,\;0.005\}$ 0.001 30 1000 1
S-agent 2 1 (128 neurons) 0.01 30 1000 1
@@ VII Evaluation Metrics
where $b_{i}$ denotes the true hidden bit in episode $i$, $\hat{b}_{i}$ is the agent’s predicted bit at episode, $\mathbb{I}(\cdot)$ is the indicator function, and $N_{\mathrm{ep}}$ is the number of evaluation episodes ($N_{\mathrm{ep}}=1000$ in our experiments).
@@ VIII-B Low-Penalty Environment
From these results, we conclude that both D-agent and S-agent outperform the C-agent, as they require fewer quantum copies to make a correct inference while achieving higher accuracy, with S-agent achieving the highest overall accuracy.. This demonstrates that hybrid agents are more resource-efficient in the low-penalty setting. This trend is further confirmed by the confusion matrices in Fig. 7, which show that both D-agent and S-agent achieve more balanced and reliable predictions across the two hidden bit values compared to the C-agent.
We note that S-agent employs a simpler quantum circuit than D-agent. In scenarios where circuit complexity is a critical consideration, S-agent offers a better trade-off by achieving higher accuracy than D-agent while requiring fewer quantum states and reduced circuit complexity, and thus lower implementation cost. This consistency highlights S-agent as a robust and resource-efficient policy across varying interaction costs.
@@ X Conclusion
Specifically, the lightweight S-agent consistently achieved reliable inference using as few as two quantum state copies, outperforming both the classical baseline and the deeper quantum policy. This finding highlights that increased quantum circuit depth does not necessarily translate to improved performance and that compact hybrid policies can offer a favorable balance between expressivity, stability, and resource efficiency. Robustness evaluations under realistic quantum noise models further showed that the hybrid agents maintain improved inference performance compared to the C-agent across a broad range of noise strengths.
@@
Another potential application for our proposed environment is information hiding. In such a setting, classical bits are encoded into quantum states using secret phase angles. The resulting quantum states are stored in quantum memory, while the corresponding angular domain is stored separately as classical side information. Information extraction is only possible for an authorized party that has access to both the stored quantum states and the associated angular domain. This knowledge enables the extractor to train an RL agent tailored to the encoding parameters, which can then interact with the quantum states to reliably recover the hidden bits. Without simultaneous access to both the quantum states and the correct angular domain, any attempt to extract the hidden information yields incorrect results, effectively concealing the embedded data. This dual-dependence on quantum memory and classical side information highlights the potential of RL-based quantum challenge–response mechanisms as a viable tool for information hiding.